OpenStudy (anonymous):

how do I solve the following equation? (1+2)^2=1^2+2(2)(1)+(2)^2

OpenStudy (skullpatrol):

There is nothing to "solve for"

OpenStudy (anonymous):

than simplify for? it says Demonstrate that your polynomial identity works on numerical relationships

OpenStudy (muzzack):

Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : \(\huge\cal\ (1+2)^2-(1^2+2*(2)*(1)+(2)^2)=0\) Step 1 : Simplify \(\huge\ 32 - 9\) Equation at the end of step 1 : 0 = 0 Step 2 : Solve 0 = 0 This equation is a tautology (Something which is always true) \(\huge\cal\ Solution: Equation~is~always~true\)

OpenStudy (skullpatrol):

$$\Huge (1+2)^2=1^2+2(2)(1)+(2)^2$$ $$\Huge(a+b)^2=a^2+2(a)(b)+(b)^2$$

OpenStudy (anonymous):

the (a+b)^2=A^2+2(a)(b)+(b)^2 is what I started with

OpenStudy (skullpatrol):

The left hand side is the square of a binomial and the right hand side is called a perfect square trinomial.

OpenStudy (skullpatrol):

Let a=1 and b=2

OpenStudy (anonymous):

Right I understand that but than it said to demonstrate with numerical relationships so I picked 1 and 2

OpenStudy (skullpatrol):

Yes, now simplify the right hand side and the left hand side then see if you get a true statement.

OpenStudy (skullpatrol):

9 = 9

OpenStudy (skullpatrol):

Is that^ true?

OpenStudy (anonymous):

true

OpenStudy (anonymous):

You get 9 for both sides

OpenStudy (skullpatrol):

Your choice of a=1 and b=2 could be changed to any numbers and would you still get a true statement?

OpenStudy (anonymous):

Yes

OpenStudy (skullpatrol):

This is called an "identity" it is a statement that is true for every value of the variable(s).

OpenStudy (anonymous):

Thank you!

OpenStudy (skullpatrol):

Thanks for asking :)